OpenStudy (anonymous):

solve log(base 27)x=1-log(base 27) (x-0.4)

OpenStudy (turingtest):

hint:\[27^{1-\log_{27}(x-0.4)}=\frac{27}{x-0.4}\]

OpenStudy (anonymous):

can you show full solution please

OpenStudy (turingtest):

No, but I can guide you through it. I have shown you the right side, what will you get by raising 27 to the power of the left side?

OpenStudy (turingtest):

\[\log_{27}x=1-\log_{27}(x-0.4)\]\[27^{\log_{27}x}=27^{1-\log_{27}(x-0.4)}\]\[x=\frac{27}{x-0.4}\]you can take it from here I hope